25. Imagine that you are buying a call option in October 2017, with a exercise price of $75. The current stock price is $82. The option price (premium) is $8.50. The option expires in January, 2018. Using this information, draw and label the price graph for this option on the day you bought it and on this graph be sure to do each of the following : (Provide your answer in the uploaded document)

a. Identify the area on the graph where the option would be in-the-money
b. Identify the area on the graph where the option would be out-of-the-money
c. Show where the price of the option would be on the day you bought it and show where the intrinsic value and the time value of the option are on the graph and calculate the intrinsic value and time value.
Answer #1

Stock Price =S = 82

Exercise Price = X = 75

In the Money Region = S > X

out-of-the-money Region = S < X  

Intrinsic Value of the option = S - X = 82 - 75 = 7

Time Value of the option = Actual option price - Intrinsic Value of the option

=8.50 - 7 = 1.50

Ans a & b :

In the Money Region & out-of-the-money Region

Call Option Price Value of Call Option Time Value of the option = $ 1.50 out-of-the-money Region Intrinsic Value of the optio

Ans c:

Time Value of the option $ 1.50

Intrinsic Value of the option $ 7